On three consecutive passes, a football team gains 6 yards, loses 24 yards, and gains 14 yards. What number represents the total net yardage? The total net yardage in yards is what? Please help. Thank you.
Answer by checkley77(12844)

6-24+14=20-24=-4 yds. ans.